Can someone please help me.
Line passing through (-10,-1) and (-8,y) is perpendicular to a line with a slope of -2/9. What is the value of y.
Thank you.

Answer by stanbon(75887)

Line passing through (-10,-1) and (-8,y) is perpendicular to a line with a slope of -2/9. What is the value of y.
---------------
Any line perpendicular to a line with slope -2/9 must have a slope
of +9/2.
---------
Slope of your line thru the given points is:
slope = [y--1]/[-8--10] = (y+1)/2
EQUATION:
(y+1)/2 = 9/2
y+1 = 9
y = 8
==============
Cheers,
Stan H.